a driver takes .20s to apply the brakes after he sees a need for it.if he is driving at a rate of 54 km/hr and the brakes are causing a deceleration of 6m/s2,find the distance travelled by the car after he sees the need to put the brakes on  ?  Innocent............

Grade:Upto college level

2 Answers

KARMA YOGESH PATEL
11 years ago

u = 54 km/hr = 15 m/s

the car will cover a distance of 300 m in the first 20 sec at const. speed of 15 m/s

after applying brakes

u = 15 m/s

a = -6 m/s2

v = 0 m/s

s = (v^2 - u^2)/2a

  = 225/12

  =18.75 m

 

hence total distance = 300 + 18.75 = 318.75 m

gaurav bansal
11 years ago

velocity of car=54km/h=15m/s

ditance covered before applying breaks=0.2*15=3m

deceleration due to breaks= 6m/s2

v2=u2+2as

0=225-2*6*s

s=18.6

total distance=18.6+3=21.6